# Bialgebras, and Lie monoid actions in Morse and Floer theory, I

### v1

@@ 2.6. Orientations and signs
Let $(-1)^{\rho^{\underline{\underline{k}}}_{\underline{\underline{l}}}}$ (resp. $(-1)^{\rho^{\underline{\underline{k}}}_{\underline{\underline{l}};i}}$) stand for the orientation of $g^{\underline{\underline{k}}}_{\underline{\underline{l}}}$ (resp. $g^{\underline{\underline{k}}}_{\underline{\underline{l}};i}$). These are the signs that will appear in ($R^{\underline{k}}_{\underline{l}}$) defining $f$-bialgebras (resp. in ($M^{\underline{k}}_{\underline{l}}$) for morphisms of $f$-algebras). In order to compute these, one first needs to understand how glueing two forests affect sign. We will need the following notations:
@@
(2.63) $v_{\geq h}(\underline{k})=(k_{i}-i^{\prime})+(k_{i+1}-1)+\cdots+(k_{a}-1).$
If $\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}$, denote
@@
Assume $\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}$ and $\underline{l}=\underline{k}^{0}\sharp\underline{k}^{1}$, then, as oriented vector spaces:
@@
For (2.65), one can go from $U^{\underline{k}^{0}}\times U^{\underline{k}^{1}}$ to $U^{\underline{k}}$ by moving successively the vectors of $U^{\underline{k}^{1}}$ corresponding to the tree $h$ in front of those of $\underline{k}^{0}$ at the right of $h$: this contributes to $(k^{1}_{h}-1)v_{\geq h}(\underline{k}^{0})$.
Identification (2.66) follows from (2.65), plus exchanging positions of $\underline{l}^{0}$ and $\underline{l}^{1}$, which adds $v(\underline{l}^{0})v(\underline{l}^{1})$.
Finally, from (2.66) and (2.65):
@@
We are now ready to determine the orientations of ${g^{\underline{\underline{k}}}_{\underline{\underline{l}}}}$ and ${g^{\underline{\underline{k}}}_{\underline{\underline{l}};f}}$.
@@
The sign exponents are given by:
${\rho^{\underline{\underline{k}}}_{\underline{\underline{l}}}}$ $=\heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}+\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}+v(\underline{l}^{0})(v(\underline{k}^{1})+v(\underline{l}^{1}))$
(2.69) $+c(\underline{k}^{1},\underline{l}^{1})(v(\underline{k}^{0})+v(\underline{l}^{0})-c(\underline{k}^{0},\underline{l}^{0})),$
(2.70) ${\rho^{\underline{\underline{k}}}_{\underline{\underline{l}};0}}$ $=\heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}+\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}+v(\underline{l}^{0})(v(\underline{k}^{1})+v(\underline{l}^{1}))+(v(\underline{k}^{0})+v(\underline{l}^{0})-c(\underline{k}^{0},\underline{l}^{0})),$
(2.71) ${\rho^{\underline{\underline{k}}}_{\underline{\underline{l}};1}}$ $=\heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}+\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}+v(\underline{l}^{0})(v(\underline{k}^{1})+v(\underline{l}^{1}))+1.$
@@
Let $h$ denote the glueing parameter in $[0,+\infty)$ and $\overrightarrow{n}=Dg^{\underline{\underline{k}}}_{\underline{\underline{l}}}.\partial_{h}$, corresponding to the outward pointing direction of $\partial^{1}\overline{K}^{\underline{\underline{k}}}_{\underline{\underline{l}}}$.
To determine $\rho^{\underline{\underline{k}}}_{\underline{\underline{l}}}$ we compare the orientation of $G^{\underline{k}}_{\underline{l}}\oplus\langle\overrightarrow{n}\rangle\oplus K^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus K^{\underline{k}^{1}}_{\underline{l}^{1}}$ with $J^{\underline{k}}_{\underline{l}}=G^{\underline{k}}_{\underline{l}}\oplus K^{\underline{k}}_{\underline{l}}$. For readability we drop signs, and then explain the sign contribution from one line to the next.
(2.72) $J^{\underline{k}}_{\underline{l}}$ $\simeq J^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus J^{\underline{k}^{1}}_{\underline{l}^{1}}$
$\simeq(G^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus K^{\underline{k}^{0}}_{\underline{l}^{0}})\oplus(G^{\underline{k}^{1}}_{\underline{l}^{1}}\oplus K^{\underline{k}^{1}}_{\underline{l}^{1}})$
$\simeq(G^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus G^{\underline{k}^{1}}_{\underline{l}^{1}})\oplus(K^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus K^{\underline{k}^{1}}_{\underline{l}^{1}})$
The signs of the isomorphism of each line are respectively $\heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}+\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}+v(\underline{l}^{0})(v(\underline{k}^{1})+v(\underline{l}^{1}))$, zero, and $c(\underline{k}^{1},\underline{l}^{1})(v(\underline{k}^{0})+v(\underline{l}^{0})-c(\underline{k}^{0},\underline{l}^{0}))$.
By assumption of being in codimension one, $c(\underline{k}^{0},\underline{l}^{0})+c(\underline{k}^{1},\underline{l}^{1})=1+c(\underline{k},\underline{l})$. Consider first the generic situation when $1=c(\underline{k}^{0},\underline{l}^{0})=c(\underline{k}^{1},\underline{l}^{1})=c(\underline{k},\underline{l})$: each groups are generated by a single vector:
(2.73) $G^{\underline{k}^{0}}_{\underline{l}^{0}}$ $=\langle v_{0}\rangle,$
(2.74) $G^{\underline{k}^{1}}_{\underline{l}^{1}}$ $=\langle v_{1}\rangle,$
(2.75) $G^{\underline{k}}_{\underline{l}}$ $=\langle v\rangle.$
Notice that $v\simeq v_{0}+v_{1}$, and modulo $v$, one has $\overrightarrow{n}=v_{1}$. Indeed, keeping the floor 0 biforest fixed, increasing $h$ corresponds to translating in the $v_{1}$ direction. Therefore,
(2.76) $G^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus G^{\underline{k}^{1}}_{\underline{l}^{1}}$ $\simeq\langle v_{0}\rangle\oplus\langle v_{1}\rangle$
$\simeq\langle v_{1}+v_{0}\rangle\oplus\langle v_{1}\rangle$
$\simeq G^{\underline{k}}_{\underline{l}}\oplus\langle\overrightarrow{n}\rangle.$
These are all oriented isomorphisms, causing no additional sign contribution.
In general, following our conventions, all free trees are at floor zero, so
(2.77) $G^{\underline{k}^{0}}_{\underline{l}^{0}}$ $=\langle v_{0}^{1},\ldots,v_{0}^{c}\rangle,$
(2.78) $G^{\underline{k}^{1}}_{\underline{l}^{1}}$ $=\langle v_{1}\rangle,$
(2.79) $G^{\underline{k}}_{\underline{l}}$ $=\langle v^{1},\ldots,v^{c}\rangle.$
Assume that the tree corresponding to $v_{1}$ faces $v_{0}^{i}$, we have:
(2.80) $(v^{1},\ldots,v^{c})$ $=(v_{0}^{1},\ldots,v_{0}^{i-1},v_{0}^{i}+v_{1},v_{0}^{i+1},\ldots,v_{0}^{c}),$
(2.81) $\overrightarrow{n}$ $=v_{1}.$
Therefore, one still has an oriented isomorphism $G^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus G^{\underline{k}^{1}}_{\underline{l}^{1}}\simeq G^{\underline{k}}_{\underline{l}}\oplus\langle\overrightarrow{n}\rangle$, leading to the same sign as in $c(\underline{k}^{0},\underline{l}^{0})=1$.
Consider now the case $GBF_{1}^{f=0}(\underline{k},\underline{l})$:
(2.82) $J^{\underline{k}}_{\underline{l}}$ $\simeq J^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus J^{\underline{k}^{1}}_{\underline{l}^{1}}$
$\simeq J^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus G^{\underline{k}^{1}}_{\underline{l}^{1}}\oplus K^{\underline{k}^{1}}_{\underline{l}^{1}}$
$\simeq G^{\underline{k}^{1}}_{\underline{l}^{1}}\oplus K^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus J^{\underline{k}^{1}}_{\underline{l}^{1}}$
The signs of the isomorphism of each line are respectively $\heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}+\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}+v(\underline{l}^{0})(v(\underline{k}^{1})+v(\underline{l}^{1}))$, zero, and $\dim G^{\underline{k}^{1}}_{\underline{l}^{1}}\dim K^{\underline{k}^{0}}_{\underline{l}^{0}}=(v(\underline{k}^{0})+v(\underline{l}^{0})-c(\underline{k}^{0},\underline{l}^{0}))$, since $c(\underline{k}^{1},\underline{l}^{1})=1$. Notice that $G^{\underline{k}^{1}}_{\underline{l}^{1}}\simeq\langle\overrightarrow{n}\rangle$.
Consider now the case $GBF_{1}^{f=1}(\underline{k},\underline{l})$:
(2.83) $J^{\underline{k}}_{\underline{l}}$ $\simeq J^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus J^{\underline{k}^{1}}_{\underline{l}^{1}}$
$\simeq G^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus K^{\underline{k}^{0}}_{\underline{l}^{0}}\oplus J^{\underline{k}^{1}}_{\underline{l}^{1}}.$
The signs of the isomorphism of each line are respectively $\heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}+\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}+v(\underline{l}^{0})(v(\underline{k}^{1})+v(\underline{l}^{1}))$, and zero. Notice that $G^{\underline{k}^{0}}_{\underline{l}^{0}}\simeq-\langle\overrightarrow{n}\rangle$, which adds an extra -1. ∎
@@ 2.7. Higher codimension corner strata
We now outline how to add the stratum of higher codimension. Suppose inductively that $\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq c-1}$ has been constructed. For $[\underline{\underline{k}},\underline{\underline{l}}]\in BF_{c}(\underline{k},\underline{l})$, in good cases, one can pick $c$ relative heights, and the glueing map will take the form
(2.84) $g^{\underline{\underline{k}}}_{\underline{\underline{l}}}\colon\left([0,+\infty]^{c}\setminus\left\{+\infty\right\}^{c}\right)\times{K}^{\underline{\underline{k}}}_{\underline{\underline{l}}}\to\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq c-1},$
along which we can glue:
(2.85) $\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq c}=\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq c-1}\cup\bigcup_{(\underline{\underline{k}},\underline{\underline{l}})\in BF_{c}(\underline{k},\underline{l})}[0,+\infty]^{c}\times{K}^{\underline{\underline{k}}}_{\underline{\underline{l}}}.$
$h_{1}$
$h_{2}$
$h_{3}$
$h_{1}$
$h_{2}$
$h_{3}$
Figure 12. Both choices of relative heights parametrize half of the corner of the same vertex of $K^{4}_{2}$.
However, as is already the case for the multiplihedron, picking the right relative heights that parametrize the whole neighborhood of the corner strata is not always possible. For example, $K^{4}_{2}\simeq J_{4}$ has a vertex with four incoming edges, see for ex. [MW10, Fig. 11]. It corresponds to the codimension 3 breaking of Figure 12, choosing the relative heights as in the picture only parametrizes half of the corner. Therefore one needs to replace $[0,+\infty]^{c}$ by a more complicated corner. Nevertheless:
Proposition 2.14.
Assuming that they have been properly constructed, $\overline{K}^{\underline{k}}_{\underline{l}}$ and $\overline{J}^{\underline{k}}_{\underline{l}}$ are compact Hausdorff spaces.
“Proof”.
Take a sequence $[\underline{U}_{n},\underline{D}_{n}]$. Up to a subsequence, assume that the type of forests are fixed (what we denoted $\varphi$ in Definition 2.1), and that all relative heights converge in $[0,+\infty]$. When some relative height equals +$\infty$+ ∞ we break the biforest, so we end up with a broken biforest which is the limit of the sequence. Furthermore the limit is unique, up to the equivalence relation involved in the definition of $BF(\underline{k},\underline{l})$. ∎
@@
Both $K=J(0)$ and $J=J(1)$ are part of a larger family $J(n)$ of $n$-grafted multiplihedra, defined as moduli spaces of biforests with $n$ grafting levels. These would correspond to sequences of group morphisms
@@
The first part corresponds to grafting levels merging together, and the second part corresponds to breaking, with the constraint that if $n_{f}=0$, then $c(\underline{k}^{f},\underline{l}^{f})=1$.
In fact, these are part of an even larger family $J(\mathcal{D})$ associated with diagrams that might not be linear, as those in [Caz23]. This will be important in defining the higher categorical structure of $f\text{-}\mathcal{B}ialg$ and $\mathcal{H}am$, we leave these questions for later.
@@ 3.2. $f$-bialgebras
A pre $f$-bialgebra is a chain complex $(A,\partial)$ with a collection of operations
@@
of degree $\deg{\alpha}^{\underline{k}}_{\underline{l}}=\dim K^{\underline{k}}_{\underline{l}}=v(\underline{k})+v(\underline{l})-c(\underline{k},\underline{l})$, satisfying the family of coherence relations:
($R^{\underline{k}}_{\underline{l}}$) $\partial({\alpha}^{\underline{k}}_{\underline{l}})=\sum_{BF_{1}(\underline{k},\underline{l})}(-1)^{\rho^{\underline{\underline{k}}}_{\underline{\underline{l}}}}\cdot{\alpha}^{\underline{k}^{0}}_{\underline{l}^{0}}\circ{\alpha}^{\underline{k}^{1}}_{\underline{l}^{1}},$
where the sum runs over $(\underline{\underline{k}},\underline{\underline{l}})=(\underline{k}^{0},\underline{l}^{0},\underline{k}^{1},\underline{l}^{1})\in BF_{1}(\underline{k},\underline{l})$, described explicitly in (2.59); and $\rho^{\underline{\underline{k}}}_{\underline{\underline{l}}}$ is given in Proposition 2.13.
A pre $f$-bialgebra is an $f$-bialgebra if furthermore it satisfies the following simplifying relations:
  • ${\alpha}^{1}_{1}=id_{A}$,
  • (Vertical forest splittings) If $\underline{l}=(1,\ldots,1)$, then (under the identifications of Section 3.1):
    ($V^{\underline{k}}_{b}$) ${\alpha}^{\underline{k}}_{\underline{l}}\simeq(-1)^{\spadesuit_{\tilde{k}}}\cdot{\alpha}^{k_{1}}_{\underline{l}}\otimes\cdots\otimes{\alpha}^{k_{a}}_{\underline{l}},$
    with $\spadesuit_{\tilde{k}}$ as in Proposition 2.6. Likewise, if $\underline{k}=(1,\ldots,1)$, then:
    ($V^{a}_{\underline{l}}$) ${\alpha}^{\underline{k}}_{\underline{l}}\simeq(-1)^{\spadesuit_{\tilde{l}}}\cdot{\alpha}^{\underline{k}}_{l_{1}}\otimes\cdots\otimes{\alpha}^{\underline{k}}_{l_{b}}.$
@@
    ($D^{\underline{k},i}_{\underline{l}}$) ${\alpha}^{\underline{k}}_{\underline{l}}\simeq{\alpha}^{\widehat{\underline{k}}}_{\underline{l}}\otimes{\alpha}^{1}_{\underline{l}}.$
@@
    ($D^{\underline{k}}_{\underline{l},i}$) ${\alpha}^{\underline{k}}_{\underline{l}}\simeq{\alpha}^{\underline{k}}_{\widehat{\underline{l}}}\otimes{\alpha}^{\underline{k}}_{1}.$
@@
Identities ($V^{\underline{k}}_{b}$), ($V^{a}_{\underline{l}}$), ($D^{\underline{k},i}_{\underline{l}}$), and ($D^{\underline{k}}_{\underline{l},i}$) come respectively from the oriented identifications of Proposition 2.6:
(3.15) ${K}^{\underline{k}}_{\underline{l}}$ $\simeq(-1)^{\spadesuit_{\tilde{k}}}\cdot{K}^{k_{1}}_{\underline{l}}\times\cdots\times{K}^{k_{a}}_{\underline{l}},$
(3.16) ${K}^{\underline{k}}_{\underline{l}}$ $\simeq(-1)^{\spadesuit_{\tilde{l}}}\cdot{K}^{\underline{k}}_{l_{1}}\times\cdots\times{K}^{\underline{k}}_{l_{b}},$
(3.17) ${K}^{\underline{k}}_{\underline{l}}$ $\simeq{K}^{\widehat{\underline{k}}}_{\underline{l}}\times{K}^{1}_{\underline{l}},$
(3.18) ${K}^{\underline{k}}_{\underline{l}}$ $\simeq{K}^{\underline{k}}_{\widehat{\underline{l}}}\times{K}^{\underline{k}}_{1}.$
@@
Let $(A,{\alpha}^{\underline{k}}_{\underline{l}})$ be an $f$-bialgebra (or more generally a pre $f$-bialgebra satisfying ${\alpha}^{1}_{1}=id$ and the ($V^{\underline{k}}_{b}$) relations). Then, with $m^{k}={\alpha}^{k}_{1}$, $(A,m^{k})$ is an $A_{\infty}$-algebra (with the sign convention below), i.e.
(3.19) $\partial(m^{k})=\sum_{\begin{subarray}{c}k+1=k^{0}+k^{1}\\ 1\leq i\leq k^{0}\end{subarray}}(-1)^{k^{0}k^{1}+i(k^{1}+1)}\cdot m^{k^{0}}\circ\left(id^{\otimes(i-1)}\otimes m^{k^{1}}\otimes id^{\otimes(k^{0}-i)}\right).$
Likewise, if $A$ is an $f$-bialgebra (or more generally a pre $f$-bialgebra satisfying ${\alpha}^{1}_{1}=id$ and the ($V^{a}_{\underline{l}}$) relations), then with $\delta^{l}=\alpha^{1}_{l}$, $(A,\delta^{l})$ is an $A_{\infty}$-coalgebra (with the sign convention below), i.e.
(3.20) $\partial(\delta^{l})=\sum_{\begin{subarray}{c}l+1=l^{0}+l^{1}\\ 1\leq j\leq l^{1}\end{subarray}}(-1)^{(l^{0}+1)j+1}\cdot\left(id^{\otimes(j-1)}\otimes\delta^{l^{0}}\otimes id^{\otimes(l^{1}-j)}\right)\circ\delta^{l^{1}}.$

### v2

@@ 2.5. Orientations and signs
(2.52) $v_{\geq h}(\underline{k})=\mathrm{Card}\left(\mathrm{Vert}(\underline{k})\cap[h,+\infty)\right)=(k_{i}-i^{\prime})+(k_{i+1}-1)+\cdots+(k_{a}-1).$
If $\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}$, recall that $\mathrm{Vert}(\underline{k})\simeq\mathrm{Vert}(\underline{k}^{0})\cup\mathrm{Vert}(\underline{k}^{1})$. Since these are ordered sets, it corresponds to a permutation of $\left\{1,\ldots,v(\underline{k})\right\}$. Observe that its signature is given by $(-1)^{\heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}}$, with
@@
Indeed, one can go from $\mathrm{Vert}(\underline{k}^{0})\cup\mathrm{Vert}(\underline{k}^{1})$ to $\mathrm{Vert}(\underline{k})$ by moving successively the vertices of $\mathrm{Vert}(\underline{k}^{1})$ corresponding to the tree $h$ in front of those of $\underline{k}^{0}$ at the right of $h$: this contributes to $(k^{1}_{h}-1)v_{\geq h}(\underline{k}^{0})$. It follows that:
@@
Assume $\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}$ and $\underline{l}=\underline{l}^{0}\sharp\underline{l}^{1}$, then, as oriented vector spaces:
@@
Equation (2.54) follows from (2.53). Identification (2.55) follows from (2.54), plus exchanging positions of $\underline{l}^{0}$ and $\underline{l}^{1}$, which adds $v(\underline{l}^{0})v(\underline{l}^{1})$. Finally, from (2.55) and (2.54):
@@
With $x^{i}=v(\underline{k}^{i})$, $y^{i}=v(\underline{l}^{i})$, $\heartsuit_{k}=\heartsuit^{\underline{k}^{1}}_{\underline{k}^{0}}$, $\heartsuit_{l}=\heartsuit^{\underline{l}^{0}}_{\underline{l}^{1}}$, the orientations of the gluing maps are given by:
(2.58) ${\rho}$ $=\heartsuit_{k}+\heartsuit_{l}+y^{0}(x^{1}+y^{1})+x^{0}+y^{0}+1,$
(2.59) ${\rho_{0}}$ $=\heartsuit_{k}+\heartsuit_{l}+y^{0}(x^{1}+y^{1})+x^{0}+y^{0}$
(2.60) ${\rho_{1}}$ $=\heartsuit_{k}+\heartsuit_{l}+y^{0}(x^{1}+y^{1})+1.$
@@
Let $K={K}^{\underline{k}}_{\underline{l}}$, $K^{i}={K}^{\underline{k}^{i}}_{\underline{l}^{i}}$, $i=0,1$; and likewise let $J={J}^{\underline{k}}_{\underline{l}}$, $J^{i}={J}^{\underline{k}^{i}}_{\underline{l}^{i}}$.
Start with $\rho_{1}$, and assume that $c(\underline{k}^{0},\underline{l}^{0})=1$. The gluing map is given by composing the isomorphisms
(2.61) $J$ $\simeq J^{0}\oplus J^{1}$
(2.62) $\simeq\mathbb{R}(-v^{0})\oplus K^{0}\oplus J^{1}.$
From (2.56), the isomorphism on the first line affects sign by $\heartsuit_{k}+\heartsuit_{l}+y^{0}(k^{1}+y^{1})$. The second line changes the sign, therefore we get $\rho_{1}$.
For $\rho_{0}$, assuming $c(\underline{k}^{1},\underline{l}^{1})=1$, The corresponding gluing map is given by:
(2.63) $J$ $\simeq J^{0}\oplus J^{1}$
(2.64) $\simeq J^{0}\oplus\mathbb{R}v^{1}\oplus K^{1}$
(2.65) $\simeq\mathbb{R}v^{1}\oplus J^{0}\oplus K^{1},$
which affects signs by $\heartsuit_{k}+\heartsuit_{l}+y^{0}(x^{1}+y^{1})+\dim J^{0}$.
For $\rho$, write
(2.66) $\mathbb{R}(v^{0}+v^{1})\oplus K$ $\simeq\mathbb{R}v^{0}\oplus K^{0}\oplus\mathbb{R}v^{1}\oplus K^{1}$
(2.67) $\simeq\mathbb{R}v^{0}\oplus\mathbb{R}v^{1}\oplus K^{0}\oplus K^{1}$
(2.68) $\simeq\mathbb{R}(v^{0}+v^{1})\oplus\mathbb{R}v^{1}\oplus K^{0}\oplus K^{1},$
which gives $\rho=\heartsuit_{k}+\heartsuit_{l}+y^{0}(x^{1}+y^{1})+\dim K^{0}+0$. ∎
@@ 2.6. Higher codimension corner strata
The partial compactifications $\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq 1}$ and $\left(\overline{J}^{\underline{k}}_{\underline{l}}\right)_{\leq 1}$ are noncompact manifolds with boundary. One can compactify them by adding inductively higher corner strata: typically, if $\underline{k}=\underline{k}^{n}\sharp\cdots\sharp\underline{k}^{0}$ and $\underline{l}=\underline{l}^{0}\sharp\cdots\sharp\underline{l}^{n}$, one should have in good cases some gluing maps of the form
(2.69) $\left((-\infty,+\infty]^{c}\setminus\left\{+\infty\right\}^{c}\right)\times{K}^{\underline{k}^{0}}_{\underline{l}^{0}}\times\cdots\times{K}^{\underline{k}^{n}}_{\underline{l}^{n}}\to\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq c-1},$
where in most cases $c=n$, and then form inductively
(2.70) $\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq c}=\left(\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq c-1}\cup\bigcup\left((-\infty,+\infty]^{c}\setminus\left\{+\infty\right\}^{c}\right)\times{K}^{\underline{k}^{0}}_{\underline{l}^{0}}\times\cdots\times{K}^{\underline{k}^{n}}_{\underline{l}^{n}}\right)/\sim.$
However, this is slightly more subtle than one might think. In particular, for some choices of multi-indices one has $c<n$, and the local corner structure might be more complicated than the standard corner $(-\infty,+\infty]^{c}$, as is already the case for the multiplihedra $J_{4}\simeq\overline{K}^{4}_{2}$, see for ex. [MW10, Fig. 11]. More details about these subtleties can be found in the first arXiv version of this paper.
@@
Both $K=J(0)$ and $J=J(1)$ are part of a larger family $J(n)$ of $n$-grafted $f$-multiplihedra, defined as moduli spaces of biforests with $n$ grafting levels. These would correspond to sequences of group morphisms
@@
The first part corresponds to grafting levels merging together, and the second part corresponds to breaking. The first author will introduce these spaces in more detail in [Cazc] in order to construct the higher categorical structure of $f\text{-}\mathcal{B}ialg$, and $\infty$-functors from the Morse complex.
@@ 3.2. $f$-bialgebras
An $f$-bialgebra is a chain complex $A$ with a collection of operations
@@
of degree $\deg{\alpha}^{\underline{k}}_{\underline{l}}=\dim K^{\underline{k}}_{\underline{l}}=v(\underline{k})+v(\underline{l})-1$, such that:
  • They satisfy the family of coherence relations:
    ($R^{\underline{k}}_{\underline{l}}$) $0=\sum_{\begin{subarray}{c}\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}\\ \underline{l}=\underline{l}^{0}\sharp\underline{l}^{1}\end{subarray}}(-1)^{\rho}\cdot{\alpha}^{\underline{k}^{0}}_{\underline{l}^{0}}\circ{\alpha}^{\underline{k}^{1}}_{\underline{l}^{1}},$
    where $\rho$ is given in Proposition 2.12. It follows from $(R^{\underline{1}_{a}}_{\underline{1}_{b}})$ that $\alpha^{\underline{1}_{a}}_{\underline{1}_{b}}$ is a differential on $(A^{b})^{a}$.
  • For $a,b\geq 1$, $\alpha^{\underline{1}_{a}}_{\underline{1}_{b}}$ coincides with the tensor product differential induced by the differential $\partial_{A}$ of $A$. We will refer to this condition as $(T^{a}_{b})$.
  • If $\underline{k}=\underline{1}_{a}$ and $\underline{l}=(1,\ldots 1,l_{j},1,\ldots,1)$, with $l_{j}\geq 2$, then
    ($V^{a}_{\underline{l}}$) ${\alpha}^{\underline{k}}_{\underline{l}}\simeq id_{(A^{j-1})^{a}}\otimes{\alpha}^{\underline{k}}_{l_{j}}\otimes id_{(A^{b-j})^{a}},$
    where $\simeq$ is in the sense of (3.4).
  • Likewise, if $\underline{l}=\underline{1}_{b}$ and $\underline{k}=(1,\ldots 1,k_{i},1,\ldots,1)$, with $k_{i}\geq 2$, then
    ($V^{\underline{k}}_{b}$) ${\alpha}^{\underline{k}}_{\underline{l}}\simeq id_{(A^{b})^{i-1}}\otimes{\alpha}^{k_{i}}_{\underline{l}}\otimes id_{(A^{b})^{a-i}},$
    where $\simeq$ is in the sense of (3.3).
  • If either $\underline{k}=\underline{1}_{a}$ and $\underline{l}$ has at least two entries $\geq 2$, or vice versa, then ${\alpha}^{\underline{k}}_{\underline{l}}=0$. We will refer to these conditions as $(V^{a}_{\underline{l}})$ and $(V^{\underline{k}}_{b})$ as well.
@@
    ($D^{\underline{k},i}_{\underline{l}}$) ${\alpha}^{\underline{k}}_{\underline{l}}\simeq{\alpha}^{\widehat{\underline{k}}}_{\underline{l}}\otimes{\widetilde{\alpha}}^{1}_{\underline{l}},$
    where ${\widetilde{\alpha}}^{1}_{\underline{l}}={\widetilde{\alpha}}^{1}_{l_{1}}\otimes\cdots\otimes{\widetilde{\alpha}}^{1}_{l_{b}}$, with ${\widetilde{\alpha}}^{1}_{1}=id_{A}$ and ${\widetilde{\alpha}}^{1}_{2}=\alpha^{1}_{2}$.
@@
    ($D^{\underline{k}}_{\underline{l},i}$) ${\alpha}^{\underline{k}}_{\underline{l}}\simeq{\alpha}^{\underline{k}}_{\widehat{\underline{l}}}\otimes\widetilde{\alpha}^{\underline{k}}_{1},$
    where ${\widetilde{\alpha}}^{\underline{k}}_{1}={\widetilde{\alpha}}^{k_{1}}_{1}\otimes\cdots\otimes{\widetilde{\alpha}}^{k_{a}}$, with ${\widetilde{\alpha}}^{1}_{1}=id_{A}$ and ${\widetilde{\alpha}}^{2}_{1}=\alpha^{2}_{1}$.
@@
The maps ${\widetilde{\alpha}}^{\underline{k}}_{\underline{l}}$ come from the spaces $\widetilde{K}^{\underline{k}}_{\underline{l}}$.
Remark 3.4.
The formula for $\rho$ in Proposition 2.12 give the orientation of the gluing map under the assumption that $c(\underline{k},\underline{l})=c(\underline{k}^{0},\underline{l}^{0})=c(\underline{k}^{1},\underline{l}^{1})=1$. Nevertheless, the expression makes sense for any $\underline{k}^{0},\underline{l}^{0},\underline{k}^{1},\underline{l}^{1}$. In particular, when $c(\underline{k}^{0},\underline{l}^{0})=0$ (resp. $c(\underline{k}^{1},\underline{l}^{1})=0$), $\rho=1$ (resp. $\rho=v(\underline{k})+v(\underline{l})+1=\deg\alpha^{\underline{k}}_{\underline{l}}\text{ mod }2$). Since we have assumed $\alpha^{\underline{k}}_{\underline{l}}=0$ for $c(\underline{k},\underline{l})\geq 2$, ($R^{\underline{k}}_{\underline{l}}$) can be written equivalently as:
(3.15) $\partial(\alpha^{\underline{k}}_{\underline{l}})=\sum_{\begin{subarray}{c}c(\underline{k}^{0},\underline{l}^{0})=1\\ c(\underline{k}^{1},\underline{l}^{1})=1\end{subarray}}(-1)^{\rho}\cdot{\alpha}^{\underline{k}^{0}}_{\underline{l}^{0}}\circ{\alpha}^{\underline{k}^{1}}_{\underline{l}^{1}},$
Where $\partial$ is the differential on $\mathrm{Hom}((A^{b})^{\left|\underline{k}\right|},(A^{\left|\underline{l}\right|})^{a})$ induced by $\partial_{A}$, and the right hand side corresponds to summing over the faces of $\partial\left(\overline{K}^{\underline{k}}_{\underline{l}}\right)_{\leq 1}$.
@@
Let $(A,{\alpha}^{\underline{k}}_{\underline{l}})$ be an $f$-bialgebra. Then, with $m^{k}={\alpha}^{k}_{1}$, $(A,m^{k})$ is an $A_{\infty}$-algebra (with the sign convention below), i.e.
(3.16) $0=\sum_{\begin{subarray}{c}k+1=k^{0}+k^{1}\\ 1\leq i\leq k^{0}\end{subarray}}(-1)^{k^{0}k^{1}+i(k^{1}+1)}\cdot m^{k^{0}}\circ\left(id^{\otimes(i-1)}\otimes m^{k^{1}}\otimes id^{\otimes(k^{0}-i)}\right).$
Likewise, if $A$ is an $f$-bialgebra, then with $\delta^{l}=\alpha^{1}_{l}$, $(A,\delta^{l})$ is an $A_{\infty}$-coalgebra (with the sign convention below), i.e.
(3.17) $0=\sum_{\begin{subarray}{c}l+1=l^{0}+l^{1}\\ 1\leq j\leq l^{1}\end{subarray}}(-1)^{(l^{0}+1)j+1}\cdot\left(id^{\otimes(j-1)}\otimes\delta^{l^{0}}\otimes id^{\otimes(l^{1}-j)}\right)\circ\delta^{l^{1}}.$
Remark 3.6.
More generally, one can show that $A^{b}$, endowed with $m^{k}=\alpha^{k}_{\underline{1}_{b}}$, is an $A_{\infty}$-algebra in the above sense. Likewise, $A^{a}$, endowed with $\delta_{l}=\tau^{a}_{l}\circ\alpha^{\underline{1}_{a}}_{l}$, is an $A_{\infty}$-coalgebra.